The sum of the ages of joy and josef now is 69. If josef is twice as old as Joe, how old were they 3 years ago?
Answer by algebra hello(55)

Let Joe's present age=x
then Josef's present age=2x
according to question,
x 2x=69
3x=69
x=23
Joe's age before 3 years= 23-3=20
Josef's age before 3 years=2*23-3=46-3=43